OpenStudy (anonymous):

parking in a student lot costs $2 for the first half hour and $1.50 for each hour thereafter. A partial hour is charged like a full hour. What is the longest time a student can park in this lot for $9.

OpenStudy (saifoo.khan):

Is that math117?

OpenStudy (saifoo.khan):

Hour 0.5: $9 - $2 = $7 remaining Hour 1.5: $7 - $1.5 = $5.5 remaining Hour 2.5: $5.5 - 1.5= $4 remaining Hour 3.5: $4 - 1.5 = $2.5 remaining HOur 4.5: $2.5 - 1.5 = $1 remaining